Treetop Corp. has a net loss (i.e., negative net income) of $3,400,000 for 2020. At January 1, 2020, Treetop had 2,500,000 shares of common stock outstanding. Treetop issued an additional 150,000 shares on April 1, and 80,000 more on July 1. On October 1, Treetop issued $5,500,000 of 4% convertible bonds at a premium of $110,000. Each $1,000 bond is convertible into 10 shares of common stock. No bonds have yet been converted. The number of shares to be used in computing diluted earnings per share for 2020 is
Answer #1

Number of shares used in Computing diluted earning per share

Weighted average shares
January 1 2500000*3/12 625000
Apr 1 2650000*3/12 662500
July 1 2730000*6/12 1365000
Total 2652500

Diluted shares = 5500000*10/1000*3/12 = 13750

Number of shares to be used in Diluted earning per share = 2652500+13750 = 2666250 shares